Solve the system by substitution and elimination method
3x+y=4
2y-8=-6x

Answer by mananth(16949)

3 x + 1 y = 4 .............1

6 x + 2 y = 8 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-6 x -2 y = -8
6 x + 2 y = 8
Add the two equations
0 x = 0
There is no solution
The graph of the lines are the same